OpenStudy (anonymous):

solve for x \[\frac{ 1 }{ \sqrt{1-x ^{2}} }=\frac{ \pi }{ 12 }\]

OpenStudy (anonymous):

\[\sqrt{1-x^{2}} = \frac{ 12 }{ \pi }\] \[1-x^{2} = \frac{ 12^{2} }{ \pi^{2} }\] \[x = \sqrt{1 - \frac{ 12^{2} }{ \pi^{2} }}\]

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

Welcome!